If the sum of the surface areas of a sphere and a cube is fixed, what is the ratio of



the radius of the sphere to the edge of the cube when the sum of their volumes is



least

1
Expert's answer
2021-11-24T14:33:25-0500

sum of the surface areas:

"S=4\pi r^2+6a^2"

edge of the cube:

"a=\sqrt{\frac{S-4\pi r^2}{6}}"


sum of volumes:

"V=4\pi r^3/3+a^3=4\pi r^3/3+(\frac{S-4\pi r^2}{6})^{3/2}"


"\frac{dV}{dr}=4\pi r^2-\frac{3}{2}\frac{8\pi r}{6}(\frac{S-4\pi r^2}{6})^{1/2}=0"


"4\pi r-\frac{3}{2}\frac{8\pi }{6}a=0"


"r/a=\frac{3}{2}\frac{8\pi }{6\cdot4\pi}=1/2"